Five square flower beds each of sides 1 m are dug on a piece of land 5 m long and 4 m wide. What is the area of the remaining part of the land?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Area of flower square bed = 1 × 1
= 1 m^2
Area of 5 square bed = 1 × 5
= 5 m^2
Area of land = 5 × 4
= 20 m^2
Remaining part of the land = Area of land–Area of 5 square bed
= 20–5
= 15 m^2
∴Remaining part of the land is 15 m^2

Thanks